Train A &B are traveling in the same direction on parallel tracks. Train A is traveling at 40mph and Train B is traveling at 44mph. Train A passes a station at 8:15am. If Train B passes the same station at 8:45am, at what time will Train B catch up to train A?
Answer by jorel1380(3719)

1/2 hour @ 40 mph=20 miles
40(n)+20=44(n)
20=4n
n=5
Train B will catch up to train A at 1:15 p.m...